### v3

@@ 4 Two Binary Program Formulations
In the following sections, we present two set-covering formulations for the problem. In Section 4.1, we extend the model proposed in Dall’Olio and Kolisch (2023) to incorporate stochastic travel times and a stochastic objective function. Besides the construction of routes, said model only decides on the profiles of teams, i.e., on lower bounds for the actual qualifications of required workers. Therefore, skill levels are implicitly aggregated, i.e., they are only considered on an aggregate level. Hence, one might obtain an integer solution that is operationally infeasible, i.e., it can not be implemented in practice. Whenever such a solution is found, we switch to an alternative formulation that considers workers based on individual skill levels. This alternative formulation, which always returns operationally feasible solutions but is considerably harder to solve, is presented in Section 4.2, alongside additional insights into operational feasibility.
@@ 4.1 Master Problem with Aggregated Workforces
The following considerations assume an underlying finite time grid given by discrete time points $\tau\in\mathcal{T}$. We first introduce general notation and several auxiliary parameters. We define a team route as a tuple $(r,q)$ consisting of a feasible route $r$ and an associated worker profile $q\in\mathcal{Q}^{r}$, which is used to execute the route. The set of team routes is denoted by $\mathcal{R}$.
Let $(r,q)\in\mathcal{R}$ be a team route. We define by $b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})\in\mathbb{N}_{\geq 0}$ the number of workers of skill level $\leq k$ that are conducting route $r$ and are occupied at time $\tau$, given worst-case travel times $t(\omega_{\mathrm{max}})$. This value is equal to $\xi_{q,k}$ for all $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and 0 else. Binary variables $\lambda^{r}_{q}$ indicate if team route $(r,q)$ is part of the solution. The problem can then be described using the following aggregated master problem, short AMP:
@@
$\underset{(r,q)\in\mathcal{R}}{\sum}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})\lambda^{r}_{q}\leq N_{k}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (8)
@@
Constraints (7) ensure that each task is part of at least one route. It is easy to see that covering a task more than once can not be part of an optimal solution. Inequalities (8) guarantee that the workforce required at any given time $\tau$ does not exceed the total available workforce on an aggregate level. The objective function (6) aims to minimize the expected finish times of tasks and incurred penalties in order to maximize buffer times. Furthermore, the AMP with only a subset of team routes $\bar{\mathcal{R}}\subset\mathcal{R}$ considered and integrality conditions (9) relaxed to $\lambda^{r}_{q}\geq 0$ is called aggregated reduced master problem (ARMP). However, for the sake of simplicity, we will still be referring to the set of columns of the ARMP as $\mathcal{R}$. We note that, unlike in standard set covering formulations, we only require the variables to be integral rather than binary. However, for any feasible instance of the AMP, there exists an optimal solution which is binary. A proof for this can be found in Appendix 9.
@@ 4.2 Master Problem with Skill-Level Specific Workforces
Let $(r,q)\in\mathcal{R}$ be a team route. As described in Section 3, $q$ is well-defined by the amount $\xi_{q,k}$ of required workers of skill level greater or equal than $k$ for all $k\in\mathcal{K}$. This information suffices to ensure that the available workforce is never exceeded on an aggregate level, but does not guarantee feasibility in terms of allocation of workers to individual tasks.
Example 4.1
Consider the following example by Dall’Olio and Kolisch (2023). There are two workers $w_{1}$ and $w_{2}$ with a skill level of $1$ and $2$, respectively, available. Assume the optimal solution consists of four tours $T_{1},T_{2},T_{3}$, and $T_{4}$ depicted in Figure 1. Tours $T_{1}$ and $T_{4}$ each require one worker with skill level $2$ while tours $T_{2}$ and $T_{3}$ each require one worker with skill level $1$. It is clear to see that the solution is feasible for the AMP. Furthermore, worker $w_{2}$ must be assigned to tours $T_{1}$ and $T_{4}$, as these require a worker with skill level $2$. Worker $w_{1}$, however, cannot be assigned to both tours $T_{2}$ and $T_{3}$ as these overlap. Therefore, while the solution selects feasible team formations for all required teams, i.e., workforce constraints are satisfied on an aggregated level, there is no feasible selection of skill compositions that does not violate workforce restrictions. Hence, the team formation, as implied by the given solution, is not implementable in practice.
Figure 1: Tour Schedule satisfying (7)–(8) (from Dall’Olio and Kolisch (2023))
@@
Example 4.2
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{q}\coloneqq\left(\xi_{q,k})\right)_{i=1,2,3}}=\left(\begin{array}[]{c}3\\ 2\\ 1\\ \end{array}\right).$
Furthermore, we assume that we have an unlimited workforce available for each skill level.
Table 2: Set of possible skill compositions for profile $q$
$s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s^{i}_{q,1}$ 1 0 0 1 0 $s^{i}_{q,2}$ 1 2 1 0 0 $s^{i}_{q,3}$ 1 1 2 2 3
Table 2 contains the set $\mathcal{S}_{q}=\left\{s_{1},s_{2},s_{3},s_{4},s_{5}\right\}$ of skill compositions that can be used to assemble a team with profile $q$. We note that each skill composition $s_{i}$ is uniquely defined by their worker usage $\left(s^{i}_{q,k}\right)_{k=1,2,3}$ per skill level. For instance, the third column in Table 2 describes skill composition $s_{2}$, which uses two workers of skill level 2 and one worker of skill level 3 to assemble a team with profile $q$. In total, team route $(r,q)$ can be used to derive five unique team routes $\left\{(r,q,s_{i}):\ i=,1\dots,5\right\}$ in $\mathcal{R}^{\mathrm{D}}$.
We use these concepts to develop a model analogous to the AMP. For each $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$, we define by $\beta_{k,\tau}^{r,s}(\omega_{\mathrm{max}})$ the number of workers of skill level $k$ required by route $r$ at time $\tau$, assuming worst-case travel times. These parameters take values $\beta_{k,\tau}^{r,s}=s_{q,k}$ for $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and $\beta_{k,\tau}^{r,s}=0$ otherwise. Using the previous definitions, we introduce the disaggregated master problem, abbreviated as DMP:
@@
$\underset{(r,q,s)\in\mathcal{R}^{\mathrm{D}}}{\sum}\beta_{k,\tau}^{r,s}(\omega_{\mathrm{max}})\lambda^{r}_{q,s}\leq N_{k}^{\mathrm{D}}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (12)
@@
Unlike in Section 4.1, constraints (12) consider workforces on an individual rather than an aggregated level. Similar to the ARMP, we call the DMP when only a subset of routes is considered and (13) is replaced by $\lambda^{r}_{q,s}\geq 0$ the aggregated reduced master problem (ARMP). For ease of reading, we will be referring to the columns of the DRMP as the set $\mathcal{R}^{\mathrm{D}}$. Moreover, we denote a solution $(\hat{\lambda}^{r}_{q})_{(r,q)\in\mathcal{R}}$ to the ARMP, which can not be extended to a solution $(\hat{\lambda}^{r}_{q,s})_{(r,q,s)\in{\mathcal{R}^{\mathrm{D}}}}$ of the DRMP as a disaggregated-infeasible solution. We note that this can be checked by solving the feasibility check, which has been proposed by Dall’Olio and Kolisch (2023) and is included in Appendix 10 of this publication. Analogous to Section 4.1, one can prove that there exists a binary solution for every feasible DMP instance.
Each disaggregated team route $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$ corresponds to one column of the DRMP. It is easy to see that $|\mathcal{R}^{\mathrm{D}}|$ is considerably larger than $|\mathcal{R}|$. A formal proof of this is provided in Appendix 13. Additionally, each column in $\mathcal{R}$ can be obtained by convex combinations of columns in $\mathcal{R}^{\mathrm{D}}$. Thus, vertices of the feasible region of the linear relaxation of the ARMP may correspond to higher-dimensional faces in the DRMP. This makes it significantly harder to cut off non-integer solutions from the feasible region; therefore, solving the DRMP using a branch-and-cut algorithm becomes substantially harder. Experimental studies have shown that solving the DRMP is, on average, around 30% slower than solving the ARMP. Therefore, we only resort to solving the DRMP once a disaggregated-infeasible solution has been identified.
@@ 5 Solution Approach
In this section, we present a Branch-Price-Cut-and-Switch algorithm to solve the problem at hand. We initially start the algorithm at the root node with a subset of columns, denoted by $\bar{\mathcal{R}}\subset\mathcal{R}$, consisting of single-task tours $(d,i,d)$ for each task $i\in\mathcal{I}$, each profile $q\in\mathcal{Q}_{i}$ and the earliest possible depot leave time $tl\coloneqq\mathrm{ES}_{i}-t_{d,i}(\omega_{\mathrm{max}})$. To ensure feasibility, we include a column that finishes each task at the latest possible time $\mathrm{LF}_{i}^{\mathrm{e}}$ and uses up the entire available workforce. We then search the branching tree using a Branch-Price-and-Cut approach. If we obtain an integer solution, we perform the feasibility check proposed by Dall’Olio and Kolisch (2023) to check if the solution is disaggregated-feasible. If the solution fails the feasibility check, we mark the current node and its sibling node as disaggregated-infeasible and restart the procedure. We note that children of marked nodes are also marked by default. Whenever a disaggregated-infeasible node is encountered with the search tree, we solve the DRMP instead of the ARMP. The selection of unexplored nodes is done via a best-first search. Table 3 extends the notation summarized in Table 1 by model- and problem-specific notation introduced throughout Sections 4 and 5.
Table 3: Model- and problem-specific notation
Sets and Variables
$\hat{\mathcal{I}}_{q}$ Set of tasks that can be executed by profile $q$ including covering
$\mathcal{R}$ Set of feasible team routes for the AMP
$\bar{\mathcal{R}}$ Subset of all feasible team routes $\mathcal{R}$
$\mathcal{R}^{\mathrm{D}}$ Set of feasible team routes for the DMP
$\mathcal{G}$ Set of indices of rank-1 CGCs present at given node
$\lambda^{r}_{q}$ Binary variable, equals to 1 if feasible team route $(r,q)$ belongs to a solution of the AMP
$\lambda^{r}_{q,s}$ Binary variable, equals to 1 if feasible team route $(r,q,s)$ belongs to a solution of the DMP
Input and Model Parameters
$t_{i,j}^{\alpha}$ largest $\alpha$-quantile of the distribution of $t_{i,j}$
$b_{k,t}^{r,q}(\omega_{\mathrm{max}})$ Number of workers with at least level $k$ required by team route $(r,q)$ at time $t$ if $\omega_{\mathrm{max}}$ occurs
$\beta_{k,t}^{r,s}$ Number of workers with skill level $k$ required by team route $(r,q,s)$ at time $t$ if $\omega_{\mathrm{max}}$ occurs
$\psi_{g}$ Value of the dual variable associated with the rank-1 CGC with index $g$
$\mathcal{G}^{q}$ Pricing network in the ARMP for profile $q$
$\mathcal{G}^{q,s}$ Pricing network in the DRMP for disaggregated profile $(q,s)$
$\mathcal{A}^{q}$ Arcs of networks $\mathcal{G}^{q}$ and $\mathcal{G}^{q,s}$
$\mathcal{V}^{q}$ Nodes of networks $\mathcal{G}^{q}$ and $\mathcal{G}^{q,s}$
$W_{i,j}(F_{i}^{P,\mathrm{max}})$ time-dynamic weight of arc $(i,j)$ depending on the worst-case finish time $F_{i}^{P,\mathrm{max}}$ of the previous task
$F_{i}^{L,\mathrm{max}}$ worst-case finish time of task $i$ in the path associated with label $L$
$u_{i}$ rank-1 CGC coefficient for task constraint associated with task $i$
$u_{k,t}$ rank-1 CGC coefficient for workforce constraint associated with skill level $k$ and time $t$
In Section 5.1, we describe the pricing problem for the ARMP. Section 5.2 focuses on the labeling algorithm used to solve the pricing problem. In Section 5.3, necessary adjustments to solve the pricing problem associated with the DRMP are explained. Section 5.4 elaborates on several acceleration strategies used to speed up the solution process. Unless otherwise stated, all considerations can be transferred to the DRMP and its set $\mathcal{R}^{\mathrm{D}}$ of disaggregated team routes.
@@ 5.1 Pricing Problem for the ARMP
$\mathbb{E}(c^{r})-\underset{i\in\mathcal{I}^{r}}{\sum}\mu_{i}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}\delta_{k,\tau}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})$
@@
We now turn our focus to the construction of the graph used to solve the ESPPRC. First, we reiterate the concept of covering profiles introduced in Dall’Olio and Kolisch (2023). For profiles $q,\bar{q}\in\mathcal{Q}$ and a set $\mathcal{K}\coloneqq\{1,\dots,K\}$ of skill levels with $K\in\mathbb{N}_{>0}$, we say that $q$ covers $\bar{q}$ if
$\xi_{\bar{q},k^{*}}$ $=\xi_{q,k^{*}}\quad\text{and}$
$\xi_{\bar{q},k}$ $\leq\xi_{q,k}\quad\forall k\in\mathcal{K}$
hold, where $k^{*}=1$ is the skill level with the lowest qualifications. For a task $i\in\mathcal{I}$, we denote by $\hat{\mathcal{Q}_{i}}$ the set of profiles covering at least one profile in $\mathcal{Q}_{i}$. It might be advantageous to execute a task $i$ using a profile $q\in\hat{\mathcal{Q}_{i}}\setminus\mathcal{Q}_{i}$, e.g., in case of a shortage of low-skilled workers. Therefore, we denote by $\hat{\mathcal{I}}_{q}\coloneqq\{i\in\mathcal{I}:q\in\hat{\mathcal{Q}_{i}}\}$ the set of tasks that might be executed using profile $q$.
@@
$\mathrm{ES}_{i}+p_{i,q}+t_{i,j}^{\alpha}>\mathrm{LF}_{j}-p_{j,q}$
holds, where
$t_{i,j}^{\alpha}\coloneqq\min\left\{t^{*}:\ \mathbb{P}(t_{i,j}\leq t^{*})\geq\alpha\right\}$
is the largest $\alpha$-quantile of the distribution of $t_{i,j}$. Therefore, we define the arc set $\mathcal{A}^{q}$ by
[…]
$\land\ \mathrm{ES}_{i}+p_{i,q}+t_{i,j}(\omega_{\max})\leq\mathrm{LF}_{j}^{\mathrm{e}}-p_{j,q}\}$
@@
$\mathrm{ES}_{j}-\mathrm{LF}_{i}^{\mathrm{e}}$ $\geq d_{i,d}{(\omega_{\mathrm{max}})}+d_{d,j}{(\omega_{\mathrm{max}})}\quad$ (14)
@@
Let $P$ be a path in $\mathcal{G}^{q}$ with depot leave time $\mathrm{tl}^{P}\geq 0$. Because the reduced costs of an arc $(i,j)\in\mathcal{A}^{q}$ are time-dependent and $\mathcal{G}^{q}$ does not contain any temporal information, arc weights are time-dynamic and specifically depend on the distribution of finish times $F_{i}^{P}$ and the worst-case finish time $F_{i}^{P}(\omega_{\mathrm{max}})$ at the previous node $i$ along $P$. For this, we set $w_{o^{\prime}}\coloneqq 0$, $\mu_{o^{\prime}}\coloneqq 0$ and $F_{o}^{P}(\omega_{\mathrm{max}})\coloneqq\mathrm{tl}^{P}-1$. Then, the weight $W_{i,j}\left(F_{i}^{P}(\omega_{\mathrm{max}})\right)$ of an arc $(i,j)\in\mathcal{A}^{q}$ is given by
$W_{i,j}\left(F_{i}^{P}(\omega_{\mathrm{max}})\right)\coloneqq w_{j}\cdot\mathbb{E}(F_{j}+P_{j}(F_{j}))-\mu_{j}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{i}^{P}(\omega_{\mathrm{max}})+1}{\overset{F_{j}^{P}(\omega_{\mathrm{max}})}{\sum}}\delta_{k,\tau}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}}).$

### v4

@@ 4 Two Binary Program Formulations
In the following sections, we present two set-covering formulations for the problem. Set-covering formulations are commonly used for vehicle routing problems in combination with column generation approaches. In such a formulation, each column represents a team performing a subset of tasks. The goal is to find a minimum cost subset of columns such that each tasks is visited (i.e., covered) exactly once, and resource or capacity constraints are satisfied. This modelling technique exploits the natural decomposable structure of routing problems. Furthermore, it allows for the consideration of the probabilistic, route-specific constraints (3) and (4) in the subproblem without the need for linearization. In Section 4.1, we extend the model proposed in dallolio_temp to incorporate stochastic travel times and a stochastic objective function. Besides the construction of routes, said model only decides on the profiles of teams, i.e., on lower bounds for the actual qualifications of required workers. Therefore, skill levels are implicitly aggregated, i.e., they are only considered on an aggregate level. Hence, one might obtain an integer solution that is operationally infeasible, i.e., it can not be implemented in practice. Whenever such a solution is found, we switch to an alternative formulation that considers workers based on individual skill levels. This alternative formulation, which always returns operationally feasible solutions but is considerably harder to solve, is presented in Section 4.2, alongside additional insights into operational feasibility. Recall that all notation used in the following is summarized in A.
@@ 4.1 Master Problem with Aggregated Workforces
The following considerations assume an underlying finite time grid given by discrete time points $\tau\in\mathcal{T}$. We first introduce general notation and several auxiliary parameters.
Let $(r,q)\in\mathcal{R}$ be a team route. We define by $b_{k,\tau}^{r,q}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\{0,\xi_{q,k}\right\}}$ the number of workers of skill level not greater than $k$ that are conducting route $r$ and are occupied at time $\tau$, given the $\gamma$-quantile scenario $\omega_{\gamma}$. This value is equal to the worker requirement $\xi_{q,k}$ of profile $q\in\mathcal{Q}$ for all $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and 0 else. Binary variables $\lambda^{r}_{q}$ indicate if team route $(r,q)$ is part of the solution. The problem can then be described using the following aggregated master problem, short AMP:
@@
$\underset{(r,q)\in\mathcal{R}}{\sum}b_{k,\tau}^{r,q}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})\lambda^{r}_{q}\leq N_{k}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (8)
@@
Constraints (7) ensure that each task is part of at least one route. It is easy to see that covering a task more than once can not be part of an optimal solution. Inequalities (8) guarantee that the workforce required at any given time $\tau$ does not exceed the total available workforce on an aggregate level. The objective function (6) aims to minimize the expected finish times of tasks and incurred penalties in order to maximize buffer times. Furthermore, the AMP with only a subset of team routes $\bar{\mathcal{R}}\subset\mathcal{R}$ considered and integrality conditions (9) relaxed to $\lambda^{r}_{q}\geq 0$ is called aggregated reduced master problem (ARMP). However, for the sake of simplicity, we will still be referring to the set of columns of the ARMP as $\mathcal{R}$. We note that, unlike in standard set-covering formulations, we only require the variables to be integral rather than binary. However, for any feasible instance of the AMP, there exists an optimal solution which is binary. A proof for this can be found in B.
@@ 4.2 Master Problem with Skill-Level Specific Workforces
Let $(r,q)\in\mathcal{R}$ be a team route. As described in Section 2, $q$ is well-defined by the amount $\xi_{q,k}$ of required workers of skill level greater or equal than $k$ for all $k\in\mathcal{K}$. This information suffices to ensure that the available workforce is never exceeded on an aggregate level, but does not guarantee feasibility in terms of allocation of workers to individual tasks. To be precise, the ARMP implicitly allows workers to migrate to another team while executing a task along their route without incurring any travel time. As this is neither realistically possible nor feasible in practical applications, this can be seen as an artifact of the AMP formulation. For an example, the reader is referred to Section 4.2 of dallolio_temp.
@@
For each $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$, we define by $\beta_{k,\tau}^{r,s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\left\{0,s_{q,k}\right\}}$ the number of workers of skill level $k$ required by route $r$ at time $\tau$, assuming scenario $\omega_{\gamma}$ occurs. These parameters take values $\beta_{k,\tau}^{r,s}=s_{q,k}$ for $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and $\beta_{k,\tau}^{r,s}=0$ otherwise. Using the previous definitions, we introduce the disaggregated master problem, abbreviated as DMP:
@@
$\underset{(r,q,s)\in\mathcal{R}^{\mathrm{D}}}{\sum}\beta_{k,\tau}^{r,s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})\lambda^{r}_{q,s}\leq N_{k}^{\mathrm{D}}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (12)
@@
Unlike in Section 4.1, constraints (12) consider workforces on an individual rather than an aggregated level. Similar to the ARMP, we call the DMP when only a subset of routes is considered and (13) is replaced by $\lambda^{r}_{q,s}\geq 0$ the aggregated reduced master problem (ARMP). For ease of reading, we will be referring to the columns of the DRMP as the set $\mathcal{R}^{\mathrm{D}}$. Moreover, we denote a solution $(\hat{\lambda}^{r}_{q})_{(r,q)\in\mathcal{R}}$ to the ARMP, which can not be extended to a solution $(\hat{\lambda}^{r}_{q,s})_{(r,q,s)\in{\mathcal{R}^{\mathrm{D}}}}$ of the DRMP as a disaggregated-infeasible solution. We note that this can be checked by solving the feasibility check, which has been proposed by dallolio_temp and is included in C of this publication. Analogous to Section 4.1, one can prove that there exists a binary solution for every feasible DMP instance.
Each disaggregated team route $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$ corresponds to one column of the DRMP. It is easy to see that $|\mathcal{R}^{\mathrm{D}}|$ is considerably larger than $|\mathcal{R}|$. A formal proof of this is provided in F. Additionally, each column in $\mathcal{R}$ can be obtained by convex combinations of columns in $\mathcal{R}^{\mathrm{D}}$. Thus, vertices of the feasible region of the linear relaxation of the ARMP may correspond to higher-dimensional faces in the DRMP. This makes it significantly harder to cut off non-integer solutions from the feasible region; therefore, solving the DRMP using a branch-and-cut algorithm becomes substantially harder. Experimental studies have shown that solving the DRMP is, on average, around 30% slower than solving the ARMP. Therefore, we only resort to solving the DRMP once a disaggregated-infeasible solution has been identified.
@@ 5 Solution Approach
In this section, we present a Branch-Price-Cut-and-Switch algorithm to solve the problem at hand. We initially start the algorithm at the root node with a subset of columns, denoted by $\bar{\mathcal{R}}\subset\mathcal{R}$, consisting of single-task tours $(d,i,d)$ for each task $i\in\mathcal{I}$, each profile $q\in\mathcal{Q}_{i}$ and the earliest possible leave time from the depot $tl\coloneqq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ES_{i}-\min\left\{t_{d,i}^{k}(\omega):k=1,\dots,|\mathcal{B}|,\ \omega\in\Omega^{k}_{d,i}\right\}}$. To ensure feasibility, we include a column that finishes each task at the latest possible time $\mathrm{LF}_{i}^{\mathrm{e}}$ and uses up the entire available workforce. We then search the branching tree using a Branch-Price-and-Cut approach. If we obtain an integer solution, we perform the feasibility check proposed by dallolio_temp to check if the solution is disaggregated-feasible. If the solution fails the feasibility check, we mark the current node and its sibling node as disaggregated-infeasible and restart the procedure. We note that children of marked nodes are also marked by default. Whenever a disaggregated-infeasible node is encountered with the search tree, we solve the DRMP instead of the ARMP. The selection of unexplored nodes is done via a best-first search.
Figure 1: Branch-Price-Cut-and-Switch
Figure 1 provides an overview over the proposed Branch-Price-Cut-and-Switch framework. In Section 5.1, we describe the pricing problem for the ARMP. Section 5.2 focuses on the labeling algorithm used to solve the pricing problem. In Section 5.3, necessary adjustments to solve the pricing problem associated with the DRMP are explained. Section 5.4 elaborates on several acceleration strategies used to speed up the solution process. Unless otherwise stated, all considerations can be transferred to the DRMP and its set $\mathcal{R}^{\mathrm{D}}$ of disaggregated team routes.
@@ 5.1 Pricing Problem for the ARMP
$\mathbb{E}(c^{r})-\underset{i\in\mathcal{I}^{r}}{\sum}\mu_{i}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}\delta_{k,\tau}b_{k,\tau}^{r,q}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$
@@
We now turn our focus to the construction of the graph used to solve the ESPPRC. For each profile $q\in\mathcal{Q}$, we define a directed graph $\mathcal{G}^{q}\coloneqq\left(\mathcal{V}^{q},\mathcal{A}^{q},b\right)$, where $b=\left(b_{k,t}^{q}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})\right)_{k\in\mathcal{K},t\in\mathcal{T}}$ are the aggregated workforce requirements of a team with profile $q$, as follows: $\mathcal{V}^{q}$ contains two nodes $o,o^{\prime}$ (called origin and destination) representing the depot and one node for each task $i\in\mathcal{I}_{q}$. Due to constraints (3)–(4), task $j$ can not be executed after task $i$ if, for any finish time of task $i$, (3) or is (4) violated. Formally, task $j$ can not be executed after task $i$ if
$t+t_{i,j}^{k(t)}(\omega_{\alpha})>LF_{j}-p_{j,q}\quad\text{or}\quad t+t_{i,j}^{k(t)}(\omega_{\mathrm{max}})>LF_{j}^{e}-p_{j,q}$ (14)
holds for all $t\in[ES_{i}+p_{i,q},LF_{i}^{e}+p_{i,q}]$, where $k(t)=\{k:t\in\mathcal{B}_{k},\mathcal{B}_{k}\in\mathcal{B}\}$ is the time bin of time period $t\in\mathcal{T}$. Condition (14) implies that, for any finish time distribution of task $i\in\mathcal{I}$, an extension along $(i,j)\in E$ would either violated the chance constraint (3) or the extended time window (4). Therefore, we define the arc set $\mathcal{A}^{q}$ by
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{for at least one }t\in[ES_{i}+p_{i,q},LF_{i}^{e}+p_{i,q}]}\}$
@@
$\mathrm{ES}_{j}-\mathrm{LF}_{i}^{\mathrm{e}}$ $\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t_{i,d}^{k}}{({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}_{d,j}}{({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})}\quad\forall k=1,\dots,|\mathcal{B}|$ (15)
@@
Let $P$ be a path in $\mathcal{G}^{q}$ with depot leave time $\mathrm{tl}^{P}\geq 0$. Because the reduced costs of an arc $(i,j)\in\mathcal{A}^{q}$ are time-dependent and $\mathcal{G}^{q}$ does not contain any temporal information, arc weights are time-dynamic and specifically depend on the distribution of finish times $F_{i}^{P}$ and the $\omega_{\gamma}$-scenario finish time $F_{i}^{P}(\omega_{\gamma})$ at the previous node $i$ along $P$. For this, we set $w_{o^{\prime}}\coloneqq 0$, $\mu_{o^{\prime}}\coloneqq 0$ and $F_{o}^{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}\coloneqq\mathrm{tl}^{P}-1$. Then, the weight $W_{i,j}\left(F_{i}^{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}\right)$ of an arc $(i,j)\in\mathcal{A}^{q}$ is given by
$W_{i,j}\left(F_{i}^{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}\right)\coloneqq w_{j}\cdot\mathbb{E}(F_{j}+P_{j}(F_{j}))-\mu_{j}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{i}^{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}+1}{\overset{F_{j}^{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}\delta_{k,\tau}b_{k,\tau}^{P,q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}.$